Solve the initial-value problem y + y - y - y = 0, y(0) = 5, y(0) = 1, y (0) = 9
Answer #1

Soin We can glll tyl-yl-y=0 y 10) = 5 y (o)= y101= 9. Welte the Ausilany Equation m3 + m² -m-1=0 m2cm+1-1 (m+10=0 im? - 1) Lputting yıcol = 1 in the above equation (1) 1 G-C2+z putting yıllo) = in the above equii) 9=City to eq from eq and ea we getputting the value above we get c=1 4 9 = 1 +₂ +4 g= 5+(2 24 y(t) = ce +Get+tget yet) lett - 4e-t t4te-t y (+) = et + 4 e-t +If you have any doubt regarding this please let me know.If you understand the solution than give me a thumbs up.thank you!!